OpenStudy (jmansedg):

Which of the points is a solution to 2x – 3y = 10? A. (-2, 2) B. (2, -2) C. (3, 2) D. (5, 1)

OpenStudy (wintersuntime):

So plug in the x and y to the equation.

OpenStudy (wintersuntime):

plug in each option to check which one gives you the answer.

OpenStudy (jmansedg):

d

OpenStudy (anonymous):

Dude, what are you doing? you know you can do his easily right? I don't want to sound mean but don't post your whole assignment on here

OpenStudy (wintersuntime):

when you plug in d 2 times 5 gives you 10 and 3 times 1 gives you 3. so 20-3=7 not 10. So d is wrong. Try again

OpenStudy (jmansedg):

b

OpenStudy (wintersuntime):

sorry i meant 10-3=7 not 20.

OpenStudy (wintersuntime):

Yes B is correct.

OpenStudy (wintersuntime):

2x2 = 4 - (-6) = 10